Solve ( x2 – 2x )( x + 3) = -2x ( x + 1)
Answer by Juhi(44)

( x2 – 2x )( x + 3) = -2x ( x + 1)
x3 + 3x2 - 2x2 -6x = -2x2 - 2x
x3 + 3x2 - 4x = 0